The 2017 balance sheet of Kerbers Tennis Shop. Inc, showed long-term debt of $5:8 million, and the 2018 balance sheet showed long-term debt of $6 million. The 2018 income statement showed an interest expense of $195,000. During 2018, the company had a cash flow to creditors of $5.000 and the cash flow to stockholders for the year was $75,000 Suppose you also know that the firms net capital spending for 2018 was $1430,000, and that the firm reduced its net working capital investment by $81.000 What was the firms 2018 operating cash flow, or OCF? (Enter your answer in dollars, not millions of dollars, e.g., 1.234.567) Operating cash flow
Answer #1

cash flow from assets=cash flow to creditors+cash flow to stockholders=-5000+75000=70000

cash flow from assets=operating cash flow- change in net working capital-net capital spending

70000=operating cash flow-(-81000)-1430000

operating cash flow=70000-81000+1430000=1419000

the above is answer..